ON COLLISION-INDUCED ABSORPTION IN PURE O_{2}, CO_{2}, AND CO_{2} - O_{2} MIXTURES.

\underline{Y. I. BARANOV}, W. J. LAFFERTY, AND G. T. FRASER, Optical Technology Division, NIST, Gaithersburg, MD 20899; A.A. VIGASIN, General Physics Institute, Russian Academy of Sciences,Vavilova 38, Moscow, 119991, Russia.

The NIST high-pressure absorption cell coupled to a DA002 Fourier-transform spectrometer has been modified in order to extend the lower temperature range of the measurements. Measurements of collision-induced absorption (CIA) of pure O_{2} and CO_{2} as well as CO_{2} - O_{2} mixtures have been recorded from room temperature down to -80^{o} C. The spectra of compressed CO_{2} in the ν_{1} - 2 ν_{2} infrared inactive Fermi dyad region consist of two anharmonically-coupled bands. Each of these bands includes a featureless CIA band on top of which is superimposed a distinctive CO_{2} dimer band. These dimer bands increase in intensity with decreasing temperature but surprisingly persist up to room temperature. Spectra of mixtures of CO_{2} and O_{2} have also been obtained. As CO_{2} is added to pure O_{2} in the absorption cell, the intensity of the O_{2} fundamental band grows rapidly. At higher CO_{2} concentrations, the band narrows and unresolved ro-vibrational structure appears at the center. Its appearance indicates that a fairly strongly bound O_{2}^{...}CO_{2} complex is formed. Attempts are presently underway to model the observed van der Waals complexes as well as the structureless CIA profiles. Binary absorption coefficients have been derived for all species studied at 0.25 cm^{−1} intervals over the range 1100 to 1800 cm^{−1}.